# Learning Adaptive and Temporally Causal Video Tokenization in a 1D Latent Space

### v1

@@ Abstract
Figure 1: AdapTok performs adaptive tokenization both temporally and across samples. Left-to-right shows token allocation adapting over time within a video, while top-to-bottom presents sample-wise allocation under different token budgets. Blue bars indicate the tokens counts used per block.
@@ 1 Introduction
The vast amount of video data available in the internet and physical world makes efficient video modeling crucial for building visual-centric agents and world models. With the rise of large language models radford2019language ; brown2020language ; touvron2023llama , auto-regressive (AR) generative modeling has become a universal paradigm for various modalities, including images and videos wang2024emu3 ; wu2024janus ; tian2024visual ; li2024synergen ; jin2024videolavit ; hong2022cogvideo .
Consequently, numerous works gupta2022maskvit ; wu2022nuwa ; hong2022cogvideo ; yan2021videogpt ; ge2022tats ; yu2023magvit ; wang2024larp have begun exploring video tokenization to quantize spatiotemporally continuous video frames into discrete token sequences, enabling autoregressive generation through causal transformers. Unlike image data, video frames exhibit temporal causality despite both being continuous signals. To better model such characteristic, some approaches yu2023magvit ; agarwal2025cosmos ; villegas2022phenaki ; wang2024omnitokenizer have incorporated temporal causal regularization, supporting online streaming processing of video frames during encoding and decoding, thereby significantly improving throughput efficiency.
However, most existing works wang2024omnitokenizer ; agarwal2025cosmos ; wang2024larp encode different frames using a fixed number of tokens, ignoring the inherent redundancy in video data. While some approaches duggal2024alit ; koike2020stochastic ; miwa2025onedpiece ; bachmann2025flextok ; shen2025cat have attempted to encode images with variable-length token sequences, few of them is extended to video data. A recent work named ElasticTok yan2024elastictok , to the best of our knowledge, made an initial attempt to encode different frames within the same video using varying numbers of tokens. However, due to its lack of global planning capabilities and constraints from 2D spatial priors, its token allocation strategy still suffers from imbalances across samples and spatial regions, resulting in suboptimal performance.
From our perspective, an efficient video tokenizer should encompass the following three key characteristics: (1) temporal causality, where the encoding and decoding of preceding frames are independent of subsequent frames, thus supporting online streaming processing; (2) 1-D latent token space, where the token allocation is decoupled from spatial structure, ensuring information density is evenly distributed in space; and (3) adaptive token allocation that flexibly adjusts token numbers based on the information of different samples under a given token budget, achieving global optimality.
@@
Extensive experiments are conducted to validate the effectiveness of our method. Notably, on UCF-101, AdapTok achieves a video reconstruction performance of FVD=28, significantly outperforming existing methods. With our adaptive token allocation strategy, AdapTok achieves Pareto optimality between performance and token count. As shown in Fig.1, AdapTok can adaptively allocate tokens based on video frame content, both temporally and sample-wise, and the reconstruction performance continues to improve as the number of tokens increases. Furthermore, we evaluated video generation performance using AdapTok as a tokenizer on UCF-101 class-conditional generation and Kinetics-600 frame prediction tasks, where our method also demonstrated significant performance improvements. Comprehensive ablation studies are also conducted to verify the necessity of each component.
@@ 2.1 Discrete Video Tokenizer
Autoregressive (AR) transformers have emerged as powerful generative models across domains, including vision. To handle continuous visual signals, VQ-VAE van2017vqvae and its extensions esser2021vqgan ; yu2021vitvqgan ; zheng2023online ; zhang2023regularized ; mentzer2023fsqvqvae introduce vector quantization to convert images into discrete code sequences. When applied to videos, tokenization methods have evolved significantly. Early approaches use per-frame tokenization gupta2022maskvit ; wu2022nuwa ; hong2022cogvideo , where image tokenizers are directly applied without temporal modeling. Block-based tokenization yan2021videogpt ; ge2022tats ; yu2023magvit improves on this by using 3D convolutions to encode short clips with local spatio-temporal context. More recently, 1-D tokenizers wang2024larp introduce learnable holistic tokens to capture global information across the entire video. Despite these advances, these methods are inherently non-causal, relying on full access to the video during encoding, which limits their applicability to streaming or autoregressive generation tasks.
To address the limitations of non-causal tokenizers, several works have explored causal temporal tokenization architectures. One line of work adopts causal convolutions for autoregressive modeling yu2023magvitv2 . Building on this, Cosmos-Tokenizer agarwal2025cosmos introduces both causal convolution layers and causal temporal attention layers for temporal modeling. In parallel, another line of work focuses on transformer-based architectures villegas2022phenaki ; wang2024omnitokenizer , where causality is enforced via attention masks, enabling autoregressive modeling using only past frames. Following this trend, our method also adopts a transformer-based causal architecture, facilitating better adaptation to downstream Physical AI applications that require a temporal causal setting.
@@ 2.2 Adaptive Tokenizer
Recent works have explored adaptive tokenization to allow variable length representations. For image tokenization, ALIT duggal2024alit dynamically adjusts the token count by recursively distilling 2D image tokens into 1D latent sequence. Another common strategy is “tail drop” koike2020stochastic , which applies higher dropout rate on the tail of the latent feature to learn ordered representations. Building on this idea, One-D-Piece miwa2025onedpiece and FlexTok bachmann2025flextok introduce tail-drop regularization for 1D image tokenizers. To support adaptive inference, CAT shen2025cat predicts image complexity using captions and large language models (LLMs), assigning each input to one of three compression ratios (8$\times$, 16$\times$, 32$\times$).
Extending adaptive tokenization to videos, ElasticTok yan2024elastictok applies tail dropping at the end of each block and supports adaptive inference by selecting token counts via a fixed threshold. However, due to the local spatial-temporal nature of these tokens—rather than forming a unified 1D sequence—it is difficult to gather informative content at the head, limiting the effectiveness of tail dropping. Moreover, relying on a fixed threshold prevents token allocation under a specified global budget. In contrast, our AdapTok adopts a 1D transformer that decouples token allocation from spatial structure. We further introduce an online scorer with IPAL for efficient token allocation, to enable globally optimal token allocation under a given token budget.
@@ 3.1.2 Block Causal Transformer
Block Causal Encoder. To design an adaptive tokenizer, it is crucial to avoid tokens that explicitly correspond to local spatial-temporal regions. To this end, the encoder employs a set of learnable latent tokens to transform the patch embeddings into a 1-D sequence. Specifically, the latent tokens $q_{\mathrm{enc}}\in\mathbb{R}^{N\times d}$ are concatenated with image tokens $e\in\mathbb{R}^{L\times d}$, forming a combined input of $(L+N)$ tokens to a block-causal transformer encoder $\mathcal{E}$. The output corresponding to the latent tokens, denoted as $\tilde{z}=\mathcal{E}(e\oplus q_{\mathrm{enc}})_{L:(L+N)}$, serves as the learned latent representation.
Across blocks, block-causal attention is applied: all tokens in a block (both latent and image tokens) can only attend to tokens from the same or previous blocks, thereby enforcing a causal dependency structure. The pattern for the encoder attention mask is shown in Fig. 2(b-i).
Block-mask Sampler. We train AdapTok to learn variable length codes over a short video block (e.g., 4 frames) by randomly dropping the tail codes for each block during training. Specifically, for each latent block $\tilde{z}_{i}$, we randomly sample the number of tokens to retain, denoted as $\ell_{i}$. After sampling $\{\ell_{i}\}_{i=1}^{K}$, a binary latent mask $m\in\{0,1\}^{M\cdot K}$ is constructed, where the first $\ell_{i}$ values in each of the $K$ blocks are set to 1, and $M$ is the total length of each block that satisfies $N=M\cdot K$.
@@
The mask $m^{\prime}$ is then applied to the latent token sequence $\tilde{z}$, where tokens with $m_{j}^{\prime}=0$ are dropped during training. The resulting sequence, denoted as $z=\mathrm{BlockTailDrop}(\tilde{z}\ |\ m^{\prime})$, is used for decoding, enabling the model to learn efficient, variable-length representations.
Quantizer. Following wang2024larp , a stochastic vector quantization (SVQ) is adopted to the quantizer $\mathcal{Q}$. Similar to VQ, a codebook $C\in\mathbb{R}^{c\times d^{\prime}}$ containing $c$ codes is maintained. Given a visual feature $z$, the cosine similarity with all code vectors in $C$ is computed, followed by a softmax to obtain a probability distribution. An index $x_{\mathrm{Ind}}$ is then sampled from this distribution via a categorical distribution:
Given the sampled index $x_{\mathrm{Ind}}$, the quantized feature $z_{q}$ is retrieved via codebook lookup, i.e., $z_{q}=C_{x_{\mathrm{Ind}}}$. To enable differentiable training, the straight-through estimator bengio2013estimating is applied.
Block Causal Decoder. The quantized tokens are then concatenated with a set of learnable image latent tokens $q_{\mathrm{dec}}\in\mathbb{R}^{L\times d}$, and the combined $(N^{\prime}+L)$ tokens are fed into a block-causal decoder $\mathcal{D}$, implemented as a transformer. The last $L$ outputs, denoted as $\hat{z}=\mathcal{D}(q_{\mathrm{dec}}\oplus z_{q},\mathcal{M}^{\prime})_{N^{\prime}:(N^{\prime}+L)}$, are used to reconstruct the video $\hat{x}\in\mathbb{R}^{T\times H\times W\times 3}$. Here, $N^{\prime}$ denotes the number of latent tokens remaining after tail-drop, and $\mathcal{M}^{\prime}$ is the attention mask applied in the decoder.This mask is derived by applying the dropout mask $m^{\prime}$ to the original block-causal attention pattern $\mathcal{M}$, as shown in Fig 2(b-ii),
@@
Within this mask, quantized latent tokens $z_{q}$ attend only to quantized tokens from the same or preceding blocks, while image latent tokens $q_{\mathrm{dec}}$ attend to all image tokens within the same block and to quantized tokens from the same or earlier blocks.
@@ 3.1.3 Training
Following previous works esser2021vqgan ; ge2022tats ; wang2024larp , the tokenizer is optimized with a composite objective:
@@
where $\mathcal{L}_{R}$ denotes the $L_{1}$ reconstruction loss, $\mathcal{L}_{VQ}$ the quantizer loss, $\mathcal{L}_{P}$ the perceptual loss zhang2018unreasonable , $\mathcal{L}_{G}$ the adversarial loss goodfellow2014generative to enhance visual quality, and $\mathcal{L}_{prior}$ the autoregressive prior loss wang2024larp for modeling the latent sequence.
@@ 3.2 Adaptive Scorer
Generating Ground-Truth Scores. At each iteration, a target block index $q\sim\mathcal{U}(0,K-1)$ is sampled. We duplicate the latent sequence ${z}$ and apply a series of latent masks $\{m^{\prime}_{p}\}$, each corresponding to a different number of tokens $\ell_{q}$, ranging from the supported range of encoding lengths. The token lengths $\ell_{i<q}$ for preceding blocks are randomly sampled, while all subsequent blocks are fully masked, i.e., $\ell_{i>q}=0$. For the $p$-th duplicated sample, the latent mask is given by:
@@
For each $m^{\prime}_{p}$, the perceptual loss $\mathcal{L}_{P}$ over the $q$-th block is computed, serving as the quality scores $s$.
@@
The mean squared error (MSE) loss $\mathcal{L}_{s}$ is computed between the predicted score $\hat{s}$ and the ground-truth quality scores $s$.
@@ 3.3 Adaptive Token allocation
1 Inputs: Video $x$, Block idx $q$;
@@ 3.4 Video Generation
Given a sequence of adaptively allocated tokens from the scorer with IPAL, an end-of-block token <EOB> is appended to the end of each block. The resulting concatenated sequence is then fed into a Llama-style transformer touvron2023llama ; touvron2023llama2openfoundation for autoregressive generation. Formally, let $y=(y_{1},y_{2},\cdots,y_{S})$ denote the multi-block adaptive sequence. The model is trained to model the probability distribution by maximize the likelihood of each token $y_{i}$ autoregressively using a standard cross-entropy loss:
@@
4 Experiments
4.1 Experimental Setups
Datasets. We train the model and evaluate its performance on two standard datasets: video reconstruction and class-conditional generation on UCF-101 soomro2012ucf101 and frame prediction on Kinetics-600 carreira2018k600 with 5-frame conditioning.
Implementation details. Following previous works ge2022tats ; yu2023magvitv2 , we adopt $16\times 128\times 128$ video clips for both training and evaluation across all tasks. Patch embeddings are extracted with a patch size of $4\times 8\times 8$, resulting in $L=1024$ tokens per video. A sequence of latent tokens $q_{\mathrm{enc}}$ is used to extract these patch embeddings, and is partitioned into $K=4$ blocks, each containing $M=512$ tokens. These latent tokens are then quantized using a codebook $C\in\mathbb{R}^{8192\times 16}$.
For adaptive token sampling, we employ a block-wise mask sampler. In each block, the number of tokens is sampled from a truncated Gaussian distribution with mean $\mu=256$, standard deviation $\sigma=128$, and bounded between $M_{\mathrm{min}}=32$ and $M_{\mathrm{max}}=512$.
AdapTok is trained on UCF-101 and Kinetics-600 for 250 epochs with a batch-size of 128. Adam kingma2014adam is adopt as the optimizer with hyperparameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$. The learning rate is linearly warmed up to $10^{-4}$ and subsequently decayed to $10^{-6}$ using a cosine scheduler.
Fréchet Video Distance (FVD) thomas2018fvd is adopt as the main evaluation metric for both reconstruction and generation. Additionally, we report MSE, PSNR and LPIPS for video reconstruction.
4.2 Main Results
Video Reconstruction. We first evaluate the video reconstruction performance of AdapTok on UCF-101, with results summarized in Tab. 2. We can observe that, even with fewer training data and latent tokens, AdapTok significantly outperforms existing causal video tokenizers, achieving an rFVD of 28 with 2048 latent tokens and 36 with 1024 tokens.
To further evaluate efficiency under varying token budgets, we compare AdapTok with several causal tokenizer baselines. In particular, we implement a strong baseline named CausalTok that shares the same block-causal architecture as AdapTok, but applies a standard non-adaptive pipeline. As shown in Fig. 3, AdapTok consistently achieves lower rFVD across different token counts. Remarkably, it achieves an FVD of 60 using only 512 tokens, outperforming most baselines, and achieves comparable performance to CausalTok while using approximately 1.8$\times$ fewer tokens.
Video Generation. For video generation, We use AdapTok to tokenize videos and train a Llama-style model to autoregressively generate token sequences. As reported in Tab. 2, our model achieves competitive performance on both class-conditional generation and frame prediction. Notably, with only 633M parameters, AdapTok-AR achieves a gFVD of 11 on Kinetics-600 and 67 on UCF-101.
Table 1: Comparison of Video Reconstruction FVD on UCF-101. All models use a causal visual tokenizer. Bold and underline indicate the best and the second best performance, respectively.
@@ 4.2 Main Results
OmniTokenizer wang2024omnitokenizer 1.4M 8,192 1,280 42
ElasticTok yan2024elastictok 356M 64,000 1,024 390
ElasticTok yan2024elastictok 356M 64,000 2,048 93
Cosmos-Tokenizer-DV agarwal2025cosmos 100M 64,000 1,280 140
AdapTok (Ours) <0.5M 8,192 512 60
AdapTok (Ours) <0.5M 8,192 1,024 36
AdapTok (Ours) <0.5M 8,192 2,048 28
@@
Phenaki villegas2022phenaki 465M 1.8B 36.4 /
MAGVIT-AR yu2023magvit 12M 306M / 265
MAGVIT-v2-AR yu2023magvitv2 1.8M 840M / 109
OmniTokenizer wang2024omnitokenizer 1.4M 650M 32.9 191
CogVideo hong2022cogvideo  $\ddagger$ 5.4M 9.4B 109.2 626
Video-LaVIT jin2024videolavit  $\ddagger$ 10M 7B / 281
TATS ge2022tats <0.5M 321M / 332
AdapTok-AR (Ours) <0.5M 633M 11 67

### v2

@@ Abstract
Figure 1: AdapTok performs adaptive tokenization both temporally and across samples. Left-to-right shows token allocation adapting over time, top-to-bottom presents sample-wise allocation under different token budgets. Blue bars indicate the tokens counts used per block (i.e., 4 frames).
@@ 1 Introduction
The vast amount of video data available in the internet and physical world makes efficient video modeling crucial for building visual-centric agents and world models. With the rise of large language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a), auto-regressive (AR) generative modeling has become a universal paradigm for various modalities, including images and videos (Wang et al., 2024c; Wu et al., 2024; Tian et al., 2024; Li et al., 2024; Jin et al., 2024; Hong et al., 2023).
Consequently, numerous works (Gupta et al., 2022; Wu et al., 2022; Hong et al., 2023; Yan et al., 2021; Ge et al., 2022; Yu et al., 2023; Wang et al., 2024a) have begun exploring video tokenization to quantize spatiotemporally continuous video frames into discrete token sequences, enabling autoregressive generation through causal transformers. Unlike image data, video frames exhibit temporal causality despite both being continuous signals. To better model such characteristic, some approaches (Yu et al., 2023; NVIDIA, 2025; Villegas et al., 2022; Wang et al., 2024b) have incorporated temporal causal regularization, supporting online streaming processing of video frames during encoding and decoding, thereby significantly improving throughput efficiency.
However, most existing works (Wang et al., 2024b; NVIDIA, 2025; Wang et al., 2024a) encode different frames using a fixed number of tokens, ignoring the inherent redundancy in video data. While some approaches (Duggal et al., 2024; Koike-Akino & Wang, 2020; Miwa et al., 2025; Bachmann et al., 2025; Shen et al., 2025) have attempted to encode images with variable-length token sequences, few of them is extended to video data. A recent work named ElasticTok (Yan et al., 2024), to the best of our knowledge, made an initial attempt to encode different frames within the same video using varying numbers of tokens. Nevertheless, due to its lack of global planning capabilities and constraints from 2D spatial priors, its token allocation strategy still suffers from imbalances across samples and spatial regions, resulting in suboptimal performance.
From our perspective, an efficient video tokenizer should encompass the following three key characteristics: (1) temporal causality, where the encoding and decoding of preceding frames are independent of subsequent frames, thus supporting online streaming processing; (2) 1-D latent token space, where the token allocation is decoupled from spatial structure, ensuring information density is evenly distributed in space; and (3) adaptive token allocation, that adjusts token numbers based on the information of different samples under a given token budget, achieving global optimality.
@@
Extensive experiments are conducted to validate the effectiveness of our method. Notably, on UCF-101, AdapTok achieves a video reconstruction performance of FVD=28, significantly outperforming existing methods. With our adaptive token allocation strategy, AdapTok achieves Pareto optimality between performance and token count. As shown in Fig.1, AdapTok can adaptively allocate tokens based on video frame content, both temporally and sample-wise, and the reconstruction performance continues to improve as the number of tokens increases. Furthermore, we evaluated video generation performance using AdapTok as a tokenizer on UCF-101 class-conditional generation and Kinetics-600 frame prediction tasks, where our method also demonstrated significant performance gains. Comprehensive ablation studies are also conducted to verify the necessity of each component.
@@ 2.1 Discrete Video Tokenizer
Autoregressive (AR) transformers have emerged as powerful generative models across domains, including vision. To handle continuous visual signals, VQ-VAE (Van Den Oord et al., 2017) and its extensions (Esser et al., 2021; Yu et al., 2022; Zheng & Vedaldi, 2023; Zhang et al., 2023; Mentzer et al., 2023) introduce vector quantization to convert images into discrete 2D code sequences. Following works such as TiTok (Yu et al., 2024b) further propose a 1D tokenizer that tokenizes images into compact 1D latent sequences. When applied to videos, early approaches use per-frame tokenization (Gupta et al., 2022; Wu et al., 2022; Hong et al., 2023), where image tokenizers are directly applied without temporal modeling. Block-based tokenization (Yan et al., 2021; Ge et al., 2022; Yu et al., 2023) improves on this by using 3D convolutions to encode short clips with local spatio-temporal context. More recently, LARP (Wang et al., 2024a) introduces holistic tokens to capture global video information. However, similar to most prior approaches, its representations remain non-causal and fixed in length, which limits their applicability to streaming or autoregressive generation tasks. In contrast, we target causal, adaptive video tokenizer that represents videos with variable‑length latent tokens, enabling online, low-latency processing while providing flexible representation granularity for downstream tasks.
To address the limitations of non-causal tokenizers, several works have explored causal temporal tokenization architectures. One line of work adopts causal convolutions for autoregressive modeling (Yu et al., 2024a). Building on this, Cosmos-Tokenizer (NVIDIA, 2025) introduces causal convolution layers and causal temporal attention layers for temporal modeling. In parallel, another line of work focuses on transformer-based architectures (Villegas et al., 2022; Wang et al., 2024b), where causality is enforced via attention masks, enabling autoregressive modeling using only past frames. Following such trend, we also adopt a transformer-based causal architecture.
@@ 2.2 Adaptive Tokenizer
[…]
Extending adaptive tokenization to videos, ElasticTok (Yan et al., 2024) applies tail dropping at the end of each block and supports adaptive inference by selecting token counts via a fixed threshold. However, due to the local spatial-temporal nature of these tokens—rather than forming a unified 1D sequence—it is difficult to gather informative content at the head, limiting the effectiveness of tail dropping. Moreover, relying on a fixed threshold prevents token allocation under a specified global budget. In contrast, our AdapTok adopts a 1D transformer that decouples token allocation from spatial structure. We further introduce an online scorer with IPAL for efficient token allocation, to enable globally optimal token allocation under a given token budget.
@@ 3.1.2 Block Causal Transformer
The block causal transformer employs a causal encoder to extract block-wise representations, a block-mask sampler to randomly sample tokens, an SVQ (Wang et al., 2024a) quantizer for discretization, and a causal decoder to reconstruct videos.
Block Causal Encoder. To design an adaptive tokenizer, it is crucial to disentangle the positional dependencies between the latent space and the local spatiotemporal regions. Motivated by Yu et al. (2024b); Wang et al. (2024a), the encoder employs learnable latent tokens to transform patch embeddings into a 1D sequence. Specifically, the latent tokens $q_{\mathrm{enc}}\in\mathbb{R}^{N\times d}$ are concatenated with video tokens $e\in\mathbb{R}^{L\times d}$, forming a combined input of $(L+N)$ tokens to a block-causal transformer encoder $\mathcal{E}$. The output corresponding to the latent tokens, denoted as $\tilde{z}=\mathcal{E}(e\oplus q_{\mathrm{enc}})_{L:(L+N)}$, serves as the latent representation.
Across blocks, block-causal attention is applied: all tokens in a block (both latent tokens $q_{\mathrm{enc}}$ and video tokens $e$) can only attend to tokens from the same or previous blocks, thereby enforcing a causal dependency structure. The pattern for the encoder attention mask is shown in Fig. 2(b-i).
Block-mask Sampler. Inspired by  Miwa et al. (2025); Bachmann et al. (2025); Yan et al. (2024), we train AdapTok to learn variable-length codes over a short video block (e.g., 4 frames) by randomly dropping the tail codes for each block during training. Specifically, for each latent block $\tilde{z}_{i}$, we randomly sample the number of tokens to retain, denoted as $\ell_{i}$. After sampling $\{\ell_{i}\}_{i=1}^{K}$, a binary latent mask $m\in\{0,1\}^{M\cdot K}$ is constructed, where the first $\ell_{i}$ values in each of the $K$ blocks are set to 1, and $M$ is the total length of each block that satisfies $N=M\cdot K$.
@@
The mask $m^{\prime}$ is then applied to the latent token sequence $\tilde{z}$, where tokens with $m_{j}^{\prime}=0$ are dropped during training. The resulting sequence, denoted as $z=\mathrm{BlockTailDrop}(\tilde{z}\ |\ m^{\prime})$, is then discretized into the quantized feature $z_{q}$.
Block Causal Decoder. The quantized tokens $z_{q}$ are concatenated with a set of learnable video latent tokens $q_{\mathrm{dec}}\in\mathbb{R}^{L\times d}$, and the combined $(N^{\prime}+L)$ tokens are fed into a block-causal decoder $\mathcal{D}$, implemented as a transformer. The last $L$ outputs, denoted as $\hat{z}=\mathcal{D}(q_{\mathrm{dec}}\oplus z_{q},\mathcal{M}^{\prime})_{N^{\prime}:(N^{\prime}+L)}$, are used to reconstruct the video. Here, $N^{\prime}$ denotes the number of quantized tokens remaining after tail-drop, and $\mathcal{M}^{\prime}$ is the attention mask applied in the decoder.This mask is derived by applying the dropout mask $m^{\prime}$ to the original block-causal attention pattern $\mathcal{M}$, as shown in Fig 2(b-ii),
@@
Within this mask, quantized tokens $z_{q}$ attend only to quantized tokens from the same or preceding blocks, while video latent tokens $q_{\mathrm{dec}}$ attend to all video tokens within the same block and to quantized tokens from the same or earlier blocks.
@@ 3.1.3 Training
Following previous works Esser et al. (2021); Ge et al. (2022); Wang et al. (2024a), the tokenizer is optimized with a composite objective:
@@
where $\mathcal{L}_{R}$ denotes the $L_{1}$ reconstruction loss, $\mathcal{L}_{VQ}$ the quantizer loss, $\mathcal{L}_{P}$ the perceptual loss (Zhang et al., 2018), $\mathcal{L}_{G}$ the adversarial loss (Goodfellow et al., 2014) to enhance visual quality, and $\mathcal{L}_{prior}$ the autoregressive prior loss (Wang et al., 2024a) for modeling the latent sequence.
@@ 3.2 Adaptive Scorer
Generating Ground-Truth Scores. At each iteration, a target block index $q\sim\mathcal{U}(0,K-1)$ is sampled. We duplicate the latent sequence z{z} and apply a series of latent masks $\{m^{\prime}_{p}\}$, each corresponding to a different number of tokens $\ell_{q}$, ranging from the supported range of encoding lengths. The token lengths $\ell_{i<q}$ for preceding blocks are randomly sampled, while all subsequent blocks are fully masked, i.e., $\ell_{i>q}=0$. For the $p$-th duplicated sample, the latent mask is given by:
@@
For each $m^{\prime}_{p}$, the perceptual loss $\mathcal{L}_{P}$ over the $q$-th block is computed, used as the quality scores $s$.
@@
The MSE loss $\mathcal{L}_{s}$ is computed between the predicted score $\hat{s}$ and the ground-truth quality scores $s$.
@@ 3.3 Adaptive Token allocation
1Inputs: Video $x$, Block idx $q$;
@@ 3.4 Video Generation
Given a sequence of adaptively allocated tokens from the scorer with IPAL, an end-of-block token <EOB> is appended to the end of each block. The resulting concatenated sequence is then fed into a Llama-style transformer (Touvron et al., 2023a; b) for autoregressive generation. Formally, let $y=(y_{1},y_{2},\cdots,y_{S})$ denote the multi-block adaptive sequence. The model is trained to model the probability distribution by maximize the likelihood of each token $y_{i}$ autoregressively using a standard cross-entropy loss:
@@
Table 1: Comparison of Video Reconstruction FVD on UCF-101. All models use a causal visual tokenizer. Bold and underline indicate the best and the second best performance, respectively. $\dagger$ denotes models trained using the same dataset and receipe as ours.
@@
OmniTokenizer (Wang et al., 2024b) 1.4M 8,192 1,280 42
ElasticTok (Yan et al., 2024) 356M 64,000 1,024 390
ElasticTok (Yan et al., 2024) 356M 64,000 2,048 93
Cosmos-Tokenizer-DV (NVIDIA, 2025) 100M 64,000 1,280 140
OmniTokenizer $\dagger$ (Wang et al., 2024b) <<0.5M 8,192 1,280 94
ElasticTok $\dagger$ (Yan et al., 2024) <<0.5M 64,000 1,022 230
CausalTok $\dagger$ <<0.5M 8,192 1,024 37
AdapTok (Ours) <<0.5M 8,192 512 60
AdapTok (Ours) <<0.5M 8,192 1,024 36
AdapTok (Ours) <<0.5M 8,192 2,048 28
@@
Phenaki (Villegas et al., 2022) 465M 1.8B 36.4 /
MAGVIT-AR (Yu et al., 2023) 12M 306M / 265
MAGVIT-v2-AR (Yu et al., 2024a) 1.8M 840M / 109
OmniTokenizer (Wang et al., 2024b) 1.4M 650M 32.9 191
CogVideo $\ddagger$ (Hong et al., 2023) 5.4M 9.4B 109.2 626
Video-LaVIT $\ddagger$ (Jin et al., 2024) 10M 7B / 281
TATS (Ge et al., 2022) <<0.5M 321M / 332
CausalTok <<0.5M 633M / 80
AdapTok-AR (Ours) <<0.5M 633M 11 67
4 Experiments
4.1 Experimental Setups
Datasets. We train the model and evaluate its performance on two standard datasets: video reconstruction and class-conditional generation on UCF-101 (Soomro et al., 2012) and frame prediction on Kinetics-600 (Carreira et al., 2018) with 5-frame conditioning.
Implementation details. Following Ge et al. (2022); Yu et al. (2024a), we adopt $16\times 128\times 128$ video clips for both training and evaluation. Patch embeddings are extracted with a patch size of $4\times 8\times 8$, resulting in $L=1024$ tokens per video. A sequence of latent tokens $q_{\mathrm{enc}}$ is used to extract these patch embeddings, and is partitioned into $K=4$ blocks, each containing $M=512$ tokens.
For adaptive token sampling, we employ a block-wise mask sampler. In each block, the token number is sampled from a truncated Gaussian distribution with mean $\mu=256$, standard deviation $\sigma=128$, and bounded between $M_{\mathrm{min}}=32$ and $M_{\mathrm{max}}=512$.
AdapTok is trained on UCF-101 and Kinetics-600 for 250 epochs with a batch-size of 128. Adam (Kingma & Ba, 2014) is adopt as the optimizer with hyperparameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$. The learning rate is linearly warmed up to $10^{-4}$ and then decayed to $10^{-6}$ using a cosine scheduler.
Fréchet Video Distance (FVD) (Unterthiner et al., 2018) is adopt as the primary metric for reconstruction and generation. Additionally, we report PSNR and LPIPS for video reconstruction.
4.2 Main Results
Video Reconstruction. We first evaluate the video reconstruction performance of AdapTok on UCF-101, with results summarized in Tab. 1. In particular, we implement a strong baseline named CausalTok that shares the same block-causal architecture as AdapTok, but applies a standard non-adaptive pipeline. To ensure fair comparison, we also reproduced previous works (i.e., ElasticTok and OmniTokenizer) using the same dataset and receipe as ours. With fewer training data and latent tokens, AdapTok significantly outperforms existing causal video tokenizers, achieving an rFVD of 28 with 2048 tokens and 36 with 1024 tokens.
[…]
Video Generation. For video generation, we use AdapTok to tokenize videos and train a Llama-style model to autoregressively generate token sequences. As reported in Tab. 2, our model achieves competitive performance on both class-conditional generation and frame prediction. With only 633M parameters, AdapTok-AR achieves a gFVD of 11 on Kinetics-600 and 67 on UCF-101. Notably, with similar reconstruction performance, AdapTok achieves better generation results than CausalTok, demonstrating the advantage of adaptively allocating tokens for different videos.
Model Scaling. To further investigate the impact of model size on performance, we trained models across 3 different sizes: AdapTok-S, AdapTok-L, and AdapTok-XL. As shown in Tab. 3, the reconstruction quality improves consistently with increasing model size. Such results indicate that AdapTok benefits from the increased model capacity, highlighting the scaling potential of our method.
Table 3: Performance of AdapTok with different model sizes. Continuous improvement is achieved as the model size increases.
Model Params rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
AdapTok-S 59M 87.32 24.23 0.151
AdapTok-L 259M 36.36 25.72 0.144
AdapTok-XL 913M 32.43 26.29 0.103